Walkenhorst Company’s machining department prepared its 2019 budget based on the following data:

Practical capacity 40,000 units
Standard machine hours per unit 2
Standard variable factory overhead $3.00 per machine hour
Budgeted fixed factory overhead $ 376,000

The department uses machine hours to apply factory overhead to production. In 2019, the department used 85,200 machine hours and incurred $637,000 in total manufacturing overhead cost to manufacture 42,010 units. Actual fixed overhead cost for the year was $386,000.

Required:

Determine for the year:

1. The fixed, variable, and total factory overhead application rates (per machine hour). (Round your answers to 2 decimal places.)

2. The total flexible budget, for factory overhead cost based on output achieved in 2019.

3. The production volume variance. State whether this variance was favorable (F) or unfavorable (U).

4. The total overhead spending variance. State whether this variance was favorable (F) or unfavorable (U).

5. The overhead efficiency variance. State whether this variance was favorable (F) or unfavorable (U).

6. The variable overhead spending variance and the fixed overhead spending variance. State whether each variance is favorable (F) or unfavorable (U).

1. fixed overhead application rate * leave blank
variable overhead application rate leave blank
total factory overhead application rate leave blank
2. total flexible budget for overhead cost leave blank
3. production volume variance **
4. total overhead efficiency variance **
5. overhead efficiency variance **
6. variable overhead spending variance **
fixed overhead spending variance **

* # answer

** enter as unfavorable or favorable
Answer #1
Answer
Ans 1 $
Variable factory overhead 3 per machine hr
Fixed overhead application rate 4.7 per machine hr
376000/40000*2
Total factory overhead application rate 7.7
ans 2
Flexible Budget for overhead cost 628060
Std machine hours allowed for actual prod= 42010*2=84020
(84020*3)+376000
ans 3
Fixed overhead production volume variance -18894
Budgeted -Applied fixed overhead 18894 F
376000- (84020*4.7)
ans 4
Tootal Overhead Spending variance 5400 U
Actual overhead-Overhead applied on actual hours
637000-(376000+(85200*3))
ans 5
Overhead efficiency variance 3540 U
(Std rate*actual hours)- Flexible Budget
(376000+(85200*3))-628060
ans 6
Variable overhead spending variance -4600
Actual Variable overhead-(Std rate*actual hours) 4600 F
(637000-386000)-(3*85200)
Fixed overhead spending variance 10000 U
Actual -Budgeted
386000-376000